# Yongyang Cai  Khyati Malik   Hyeseon Shin

### v2

@@ 2.3 Economic System
where $c_{i,t}$ is per capita consumption and $\gamma$ is the inverse of intertemporal elasticity of substitution. We follow DICE-2016R Nordhaus (2017) and set $\gamma=1.45$.
@@
where $K_{i,t}$ is capital stock, $L_{i,t}$ is labor, and $\alpha=0.3$ (following Nordhaus (2017)) is the elasticity of gross output with respect to capital. The total factor productivity $A_{i,t}$ is assumed to follow an exogenous and deterministic time-varying trend, which is calibrated with the no-climate-impact GDP projections in Burke et al. (2018). Labor $L_{i,t}$ is also assumed to follow the exogenous population path of region $i$ from the SSP2 scenario (Samir and Lutz, 2017).
@@
where $T_{t}$ is the global average temperature increase in degrees Celsius above the preindustrial level. We calibrate the climate damage parameters based on the projections of GDP loss in Kahn et al. (2021). We follow DICE and assume emission abatement cost is
$\Phi_{i,t}=b_{1,i,t}\mu_{i,t}^{b_{2,i}}Q_{i,t},$
where $b_{1,i,t}=(b_{1,i}+b_{3,i}\exp(-b_{4,i}t))\sigma_{i,t}$. We calibrate the abatement cost parameters, $b_{1,i}$, $b_{2,i}$, $b_{3,i}$, and $b_{4,i}$, by exploiting the projections under different carbon tax scenarios in Ueckerdt et al. (2019), from which we recover the marginal abatement costs and calibrate the parameters we need.
@@
where $I_{i,t}$ is investment and $m_{t}$ is the market price of carbon emission permits. The cost of purchasing emission permits from other regions results in reduced consumption, while selling emission permits increases consumption. The capital stock evolves according to the standard neoclassical growth model:
@@ 3 Solving for the Equilibrium
Based on the model components introduced in Section 2, we now define equilibrium under cooperative and noncooperative models in the ETS. These two extreme cases place bounds on what we would observe in a real world scenario.
3.1 The Cooperative and the Noncooperative Model
In the cooperative model, all regions work together to maximize their combined welfare. More specifically, the objective function is to maximize the overall sum of discounted social utility across all regions as follows:
${\color[rgb]{0,0,0}\begin{split}\underset{c_{i,t},E_{i,t}^{P},\mu_{i,t}}{\max}&\sum_{t=0}^{\infty}\sum_{i\in\mathcal{I}}\beta^{t}u(c_{i,t})L_{i,t}\end{split},}$ (13)
where $\beta$ is the discount factor. We follow DICE-2016R (Nordhaus, 2017) and set $\beta=0.985$. Note that under cooperation there is no market equilibrium permit price, so we assume permit prices follow an exogenous path. Therefore, we define the cooperative model equilibrium given an exogenous path of permit prices.
DEFINITION:
Given the initial capital and cumulative global emissions, $\{K_{i,0},\mathcal{E}_{0}:\ i\in\mathcal{I}\}$, and the exogenous paths of emission caps and permit prices, $\{\overline{E}_{i,t},m_{t}:\ i\in\mathcal{I},t\geq 0\}$, the dynamic equilibrium for the cooperative model is a sequence of quantities $\{c_{i,t},E_{i,t}^{P},\mu_{i,t},K_{i,t},\mathcal{E}_{t},T_{t}:\ i\in\mathcal{I},t\geq 0\}$ that solve the maximization problem (13) subject to equations (7)-(12).
In the noncooperative model, individual regions maximize their own lifetime social welfare, which represents another extreme but relatively more realistic perspective. The maximization problem for each region $i$ is defined as
$\begin{split}\underset{c_{i,t},E_{i,t}^{P},\mu_{i,t}}{\max}&\sum_{t=0}^{\infty}\beta^{t}u(c_{i,t})L_{i,t}\end{split}.$ (14)
@@ 3.2 The Algorithm for the Noncooperative Model
Here we outline the algorithm we develop to obtain the optimal solution for our model. Since $\gamma=1.45$ implies utilities of per capita consumption have negative values and that long-run per capita consumption will be large, the long-run per-capita utility has a small magnitude. Together with the discount factor $\beta=0.985$, the discounted utilities after 300 years are nearly zero and have little impact on the solution in the first 100 years. Thus, we can transform the infinite horizon models to finite horizon models, where region $i$’s social welfare is rewritten as
$\sum_{t=0}^{299}\beta^{t}u(c_{i,t})L_{i,t}+\beta^{300}u\left(\frac{0.75Y_{i,300}}{L_{i,300}}\right)\frac{L_{i,300}}{1-\beta}.$
Here the second term is an approximation of the sum of discounted utility from time $t=300$ to infinity, assuming that consumption at any $t\geq 300$ is identical to 75 percent of the output at $t=300$ and that the exogenous population after 300 years stays at its value at $t=300$. The algorithm to solve the noncooperative model is as follows:
@@
1. Initialization. Give an initial guess of permit prices $\{m_{t}^{0}:\ t\geq 0\}$ and solve the cooperative model with emission cap and trade. The optimal net emissions and the amount of emission permits traded with the other regions are denoted as $\{E_{i,t}^{0},E_{i,t}^{P,0}:\ t\geq 0\}$, for each region $i$. Iterate through steps 2, 3 and 4 for $j=0,1,...$, until convergence.
@@
2. Maximization Step. Solve the maximization problem (14) without the market clearing condition (12) for each region. Given the permit prices $\{m_{t}^{j}:\ t\geq 0\}$, we solve the optimization problem (14) without the market clearing condition (12) for region 1, where we fix the other regions’ net emissions at $\{E_{i,t}^{j}:\ i=2,...,12,\ t\geq 0\}$. We then obtain region 1’s optimal solution, where the optimal net emissions and the amount of emission permits traded with the other regions are denoted as $\{E_{1,t}^{*},E_{1,t}^{P,j+1}:\ t\geq 0\}$ Similarly, we obtain $\{E_{i,t}^{*},E_{i,t}^{P,j+1}:\ t\geq 0\}$ for regions $i=2,...,12$.
@@
3. Update the emission permit prices and net emissions. After solving for the optimization problem of all regions, update the permit prices and net emissions as
$m_{t}^{j+1}$ $=$ $m_{t}^{j}\exp\left(\omega\sum_{i\in\mathcal{I}}E_{i,t}^{P,j+1}\right),$
$E_{i,t}^{j+1}$ $=$ $\omega E_{i,t}^{*}+(1-\omega)E_{i,t}^{j},\ \ \forall i\in\mathcal{I},$
where $\omega=0.1$ is a weight parameter and $\sum_{i\in\mathcal{I}}E_{i,t}^{P,j+1}$ is the net quantity of traded emission permits.
@@
4. Check the convergence criterion. Check if $m_{t}^{j+1}\simeq m_{t}^{j}$, $E_{i,t}^{j+1}\simeq E_{i,t}^{j}$, and $E_{i,t}^{P,j+1}\simeq E_{i,t}^{P,j}$ for every region $i$ and $t\geq 0$. If so, stop the iteration. Note that $m_{t}^{j+1}=m_{t}^{j}$ implies that the market clearing condition (12) holds at the solution.
@@
This algorithm embodies the concept of market equilibrium. When there is a positive net quantity of traded emission permits in the market, indicating excess demand, we increase the permit prices. Conversely, when there is a negative net quantity of traded emission permits, indicating excess supply, we lower the permit prices. This mechanism ensures that the market reaches a balance between supply and demand, and thus the market clears. Furthermore, this algorithm guarantees that each region obtains its optimal solution and reaches an equilibrium state. In other words, no region has the incentive to deviate from the Nash equilibrium for the noncooperative model solution. Note that the equilibrium of the noncooperative model is an open-loop Nash equilibrium (OLNE), which provides a solution path over time depending on the initial state. The concept of the OLNE assumes the regions enter an agreement to commit to the solution path over time at the initial period. While this type of equilibrium concept might not be as satisfactory as the feedback Nash equilibrium concept, it has the computational advantages of solving open-loop versus feedback, particularly when the dimension of the state space is large and there are occasionally binding constraints as in our case.
4 SCC, MAC, and the Optimal Carbon Tax
In this section, we first define the regional social cost of carbon (SCC) and marginal abatement cost (MAC) and explain how the optimal carbon tax can be introduced together with the global ETS, in relation to the SCC and MAC.
4.1 SCC and MAC
The social cost of carbon (SCC), a central concept in the climate change literature, is widely used to quantify the monetary value of climate damages induced by an additional unit of carbon emissions. While the SCC is often calculated in a global social planner’s problem (e.g., the DICE model), we consider the SCC for each region. Similar to van der Ploeg and de Zeeuw (2016) and Cai et al. (2023), we define the cooperative SCC of a region as the marginal rate of substitution between global emissions and regional capital as follows:
$\mathrm{SCC}_{i,t}^{\mathrm{Coop}}=\frac{-1,000(\partial V_{t}^{\mathrm{Coop}}/\partial\mathcal{E}_{t})}{\partial V_{t}^{\mathrm{Coop}}/\partial K_{i,t}},$ (15)
for region $i$, where $V_{t}^{\mathrm{Coop}}$ is the value function of the cooperative model at time $t$, depending on the state variables $\{K_{i,t},\mathcal{E}_{t}:\ i\in\mathcal{I}\}$; that is,
$V_{t}^{\mathrm{Coop}}(K_{1,t},...,K_{12,t},\mathcal{E}_{t})=\max_{c_{i,s},E_{i,s}^{P},\mu_{i,s}}\sum_{s=t}^{\infty}\beta^{t}\sum_{i=1}^{12}u(c_{i,s})L_{i,s}.$
Similarly, the noncooperative SCC of a region is defined as
$\mathrm{SCC}_{i,t}^{\mathrm{OLNE}}=\frac{-1,000(\partial V_{i,t}^{\mathrm{OLNE}}/\partial\mathcal{E}_{t})}{\partial V_{i,t}^{\mathrm{OLNE}}/\partial K_{i,t}},$ (16)

### v3

@@ 3.3 Economic System
where $c_{i,t}$ is per capita consumption and $\gamma$ is the inverse of intertemporal elasticity of substitution. We follow DICE-2016R Nordhaus, (2017) and set $\gamma=1.45$.
@@
where $K_{i,t}$ is capital stock, $L_{i,t}$ is labor, and $\alpha=0.3$ (following Nordhaus, (2017)) is the elasticity of gross output with respect to capital. The total factor productivity $A_{i,t}$ is assumed to follow an exogenous and deterministic time-varying trend, which is calibrated with the GDP projections in Burke et al., (2018). Labor $L_{i,t}$ is also assumed to follow the exogenous population path of region $i$ from the SSP2 scenario (Samir and Lutz,, 2017).
@@
where $T_{t}$ is the global average temperature increase in °C above the pre-industrial level. We calibrate the climate damage parameters based on the projections of GDP in Burke et al., (2018). We follow DICE and assume that the emission abatement cost is
$\Phi_{i,t}=b_{1,i,t}\mu_{i,t}^{b_{2,i}}Q_{i,t},$ (18)
where $b_{1,i,t}=(b_{1,i}+b_{3,i}\exp(-b_{4,i}t))\sigma_{i,t}$. We calibrate the abatement cost parameters, $b_{1,i}$, $b_{2,i}$, $b_{3,i}$, and $b_{4,i}$, by exploiting the projections under different carbon tax scenarios in Ueckerdt et al., (2019), from which we recover the marginal abatement costs and calibrate the parameters we need.
@@
where $I_{i,t}$ is investment and $m_{t}$ is the market price of carbon emission permits. The cost of purchasing emission permits from other regions results in reduced consumption, while selling emission permits increases consumption, but this impact is little as the cost or profit from permit trading is small relative to the consumption or output. The capital stock evolves according to the standard neoclassical growth model:
@@ 4 Solving for the Equilibrium
Building on the model components outlined in Section 3, we now define the noncooperative equilibrium of our multi-region dynamic model with the ETS and present an algorithm used to obtain the Nash Equilibrium solution.
4.1 The Noncooperative Equilibrium
In the noncooperative model, the regional social planner of each region maximizes the region’s own lifetime social welfare. The maximization problem for each region $i$ is defined as
$\max_{c_{i,t},E_{i,t}^{P},\mu_{i,t}}\sum_{t=0}^{\infty}\beta^{t}u(c_{i,t})L_{i,t},$ (22)
where $\beta$ is the discount factor. We follow DICE-2016R (Nordhaus,, 2017) and set $\beta=0.985$. Since one region’s emissions will influence the global average temperature, and therefore other regions’ output, the maximization problems of all regions have to be solved simulatenously as a dynamic game. Then we define the dynamic Nash equilibrium of the economy as follows.
@@ 4.1 The Noncooperative Equilibrium
Note that the equilibrium concept in our noncooperative model is an Open-Loop Nash equilibrium (OLNE), which provides a solution path over time depending on the initial state. In an OLNE, regions commit to the strategies over time for their decision variables—consumption ($c_{i,t}$), emission purchase ($E_{i,t}^{P}$), and emission control rate ($\mu_{i,t}$)— at the initial period and cannot change their behavior over time. This concept contrasts with the Markov Perfect Equilibrium (MPE), where regions may make multiple decisions over time, allowing adaptation in their strategies. While the OLNE concept may be less satisfactory than the MPE concept (since OLNE is not subgame perfect), it has the computational advantages of solving open-loop versus feedback, particularly when the dimension of the state space is large and there are occasionally binding constraints as in our case.
@@ 4.2 The Algorithm for the Noncooperative Model
Here we outline the algorithm we develop to obtain the optimal solution for our model. Since $\gamma=1.45$ implies utilities of per capita consumption have negative values and that long-run per capita consumption will be large, the long-run per-capita utility has a small magnitude. Together with the discount factor $\beta=0.985$, the discounted utilities after 300 years are nearly zero and have little impact on the solution in the first 100 years. Let
$V_{i,300}(K_{1,300},...,K_{12,300},\mathcal{E}_{300})=u\left(\frac{0.75Y_{i,300}}{L_{i,300}}\right)\frac{L_{i,300}}{1-\beta}$
be a terminal value function at the terminal year 300, which approximates the present value of utilities after 300 years, assuming that consumption at any $t\geq 300$ is equal to 75 percent of the output $Y_{i,300}$ at $t=300$ and that the exogenous population after 300 years stays at its value at $t=300$. Note that $Y_{i,300}$ is computed with a function of the terminal state $K_{i,300}$ and $\mathcal{E}_{300}$. Thus, we can transform the infinite horizon models to finite horizon models, where region $i$’s social welfare is rewritten as
$\sum_{t=0}^{299}\beta^{t}u(c_{i,t})L_{i,t}+\beta^{300}V_{i,300}(K_{1,300},...,K_{12,300},\mathcal{E}_{300}),$
We also numerically verify that this time horizon truncation at 300 years has little impact on the OLNE solution in the first 100 years, by solving the same model but with a time horizon truncation at 400 years. This time horizon truncation method is common in solving infinite-horizon non-stationary models. For example, DICE-2016 truncates its model’s infinite horizon to 500 years, with a terminal value function being zero everywhere, for obtaining its numerical solution.
The algorithm to solve the noncooperative model is as follows:
@@
1. Initialization. Set an initial guess of permit prices $\{m_{t}^{0}:\ t\geq 0\}$, and emissions $\{E_{i,t}^{0}:\ t\geq 0\}$. Iterate through steps 2, 3 and 4 for $j=1,2,...$, until convergence.
@@
2. Maximization Step at iteration $j$. For each region $i$, solve the maximization problem (22) without the market clearing condition (21), assuming the permit prices $\{m_{t}^{j-1}:\ t\geq 0\}$ and other regions’ emissions $\{E_{i^{\prime},t}^{j-1}:\ i^{\prime}\neq i,\ t\geq 0\}$ are given from the initialization step when $j=1$ or Step 3 at iteration $j-1$ when $j>1$. The optimal emissions and permits purchased for region $i$ are denoted $\{E_{i,t}^{*,j},E_{i,t}^{P,j}:\ t\geq 0\}$.
@@
3. Update Step at iteration $j$. After solving for the optimization problem of all regions respectively in Step 2, update the permit prices and emissions as
$m_{t}^{j}$ $=$ $m_{t}^{j-1}\exp\left(\omega\sum_{i\in\mathcal{I}}E_{i,t}^{P,j}\right),$
$E_{i,t}^{j}$ $=$ $\omega E_{i,t}^{*,j}+(1-\omega)E_{i,t}^{j-1},\ \ \forall i\in\mathcal{I},$
where $\omega=0.1$ is a weight parameter and $\sum_{i\in\mathcal{I}}E_{i,t}^{P,j}$ is the net quantity of traded emission permits.
@@
4. Check the convergence criterion. Check if $m_{t}^{j}\simeq m_{t}^{j-1}$, $E_{i,t}^{j}\simeq E_{i,t}^{j-1}$, and $E_{i,t}^{P,j}\simeq E_{i,t}^{P,j-1}$ for every region $i$ and $t\geq 0$. If so, stop the iteration, otherwise go to Step 2 by increasing $j$ with 1. Note that $m_{t}^{j}=m_{t}^{j-1}$ implies that the market clearing condition (21) holds at the solution.
@@
This algorithm embodies the concept of market equilibrium. When there is a net positive quantity of traded emission permits in the market, indicating excess demand, we increase the permit prices. Conversely, when there is a net negative quantity of traded emission permits, indicating excess supply, we lower the permit prices. This mechanism ensures that the market reaches a balance between supply and demand, and thus the market clears. Furthermore, this algorithm guarantees that each region obtains its optimal solution and reaches an equilibrium state. In other words, no region has the incentive to deviate from the Nash equilibrium for the noncooperative model solution.
5 Data and Calibration
To calibrate our model with data, we have two objectives. The first objective is to generate regional emission cap pathways ($\overline{E}_{i,t}$) for future periods, constraining the constituent nations of each region to meet their emissions commitments under the Glasgow Pact and the long-term net zero emission targets. The second objective is to determine parameters for total factor productivity ($A_{i,t}$), carbon intensity ($\sigma_{i,t}$), abatement cost ($b_{1,i}$, $b_{2,i}$, $b_{3,i}$, $b_{4,i}$), and climate damage ($\pi_{1,i},\pi_{2,i}$) that reflect the future projections provided in recent studies (Burke et al.,, 2018; Ueckerdt et al.,, 2019; Kahn et al.,, 2021). We obtain these region-specific parameters, which capture regional heterogeneities in GDP growth, emissions, technologies, and climate damages.
Our model’s 12 regions are formed by aggregating 152 countries around the world, following the regional classification in the RICE model (Nordhaus and Yang,, 1996; Nordhaus, 2010b, ). For the initial year 2020, we obtain country-level historical data on population (billions), capital (\$ trillions, 2020), GDP (\$ trillions, 2020), and emissions (GtC) from the World Bank. For future projections on population growth, we use the SSP2 scenario (Samir and Lutz,, 2017).
5.1 Regional Emission Cap Pathways
Since there is no global cap-and-trade system or ETS currently in place, we consider the emission commitments outlined in the Paris Agreement and the COP26 Glasgow Climate Pact. Under the Paris Agreement, 195 countries or regions set Nationally Determined Contributions (NDCs) that specify their near-term targets for 2025 or 2030, along with long-term net-zero commitments for 2050 to 2070. These near-term targets were further strengthened during the Glasgow Climate Pact in 2021, with most countries and regions submitting updated or new NDCs. [^10] Individual NDC documents were obtained from the following source: UNFCCC NDC Registry (https://unfccc.int/NDCREG) We collect reports of the most updated Nationally Determined Contributions (NDCs) after the Glasgow Climate Pact and obtain the target years to reach net zero emissions of different countries from Climate Action Tracker. [^11] See https://climateactiontracker.org/. Based on these datasets, we create the baseline regional emission cap pathways for future periods following the strategy detailed below.
As the first step, we obtain the near-term emission targets for each country. In their NDC reports, most countries express targets as a specific percentage reduction in emissions by 2030 (or, in some cases, 2025) compared to their Business As Usual (BAU) emission level at some base year. Some countries, including China, Chile, Malaysia, Singapore, and Tunisia specified their targets as a percentage reduction in carbon intensity instead of a percentage reduction in emissions. For the countries that did not make a specific emissions reduction pledge, we assume that their carbon intensity reduction and emission reduction percentages are the same as those of the most populous country in that region.
Next, we generate annual emission cap pathways based on the regions’ historical emission levels (World Bank,, 2020), their emission targets for 2030 (or 2025), and net zero emission target years. We use five-year emissions data (2014 - 2018) from the World Bank and the INDC emission targets in 2030 (or 2025) to fit a quadratic function and use this fitted function to project the emission pathways for the periods between 2018 and 2030. Emission projections for the years between 2030 (or 2025) and the net zero emission target year are obtained by linearly interpolating the emissions. After obtaining the country-level emission pathways, we aggregate them to find the regional emission cap pathways. [^12] We also find that our aggregated regional emission cap pathways are close to those used in Nordhaus, 2010b . To assess the implications of varying stringency in the emission caps, we also create additional emission cap pathways by choosing alternative net zero scenarios, wherein we assume that all regions achieve net zero emissions by 2050 (most stringent), 2070, or 2090 (most lax). See Figures A.5 and A.6 in Appendix A.3.4 for the regional emission cap pathways in the baseline cap scenario and the different emission cap pathways at the global level.
5.2 Total Factor Productivity and Climate Damage
We follow Cai et al., (2023) to calibrate total factor productivity (TFP), $A_{i,t}$, and the climate damage parameters $\pi_{1,i}$ and $\pi_{2,i}$, based on Burke et al., (2018), who provide the projected GDP of 165 countries till 2099 assuming no climate-related impacts, and their GDP assuming the contemporaneous climate impact of the RCP4.5 temperature $T_{t}^{\mathrm{RCP4.5}}$ [^13] We choose the RCP4.5 scenario instead of the other RCP scenarios, because the RCP4.5 scenario is the closest which covers the range of temperature in our solution. until 2049, under the SSP2 population pathway. [^14] It is nontrivial to use historical data to calibrate future TFP, particularly when we need to isolate the climate impacts from the data. For simplicity, we use the projected GDP for the calibration in this paper. We also use the TFP growth values in RICE to do sensitivity analysis, and find our results are still robust (see Appendix A.4.6). We aggregate these projections according to our 12 regions and employ the SSP2 population scenario to obtain regional GDP per capita estimates, $y_{i,t}^{\mathrm{BDD,NoCC}}$ under no climate impact and $y_{i,t}^{\mathrm{BDD}}$ under climate impact, which are used to calibrate the regional TFP, $A_{i,t}$, and the climate damage parameters, $\pi_{1,i}$ and $\pi_{2,i}$. For each region $i$, the dynamic path of the TFP is modeled by the relationship $A_{i,t+1}=A_{i,t}\exp\left(g_{i,t}\right)$, where $g_{i,t}$ is the growth rate of $A_{i,t}$ at time $t$. When $t<80$ (i.e., within this century), we assume
$g_{i,t}=g_{i,0}\exp\left(-d_{i}t\right).$ (23)
For $t\geq 80$ (i.e., beyond this century), since the cumulative effect is huge for a long horizon, it is often inappropriate to simply extrapolate TFP growth rate using the formula (23). Therefore, we follow RICE to generate $g_{t,i}$ for $t\geq 80$. See Appendix A.3.5 for the details.
In our structural estimation, we obtain $(g_{i,0},d_{i},\pi_{1,i},\pi_{2,i})$ by solving the following minimization problem:
$\min_{g_{i,0},d_{i},\pi_{1,i},\pi_{2,i}}\ \sum_{t=0}^{79}\left(\frac{y_{i,t}^{\mathrm{NoCC}}}{y_{i,0}^{\mathrm{NoCC}}}-\frac{y_{i,t}^{\mathrm{BDD,NoCC}}}{y_{i,0}^{\mathrm{BDD,NoCC}}}\right)^{2}+\sum_{t=0}^{29}\left(\frac{y_{i,t}}{y_{i,0}}-\frac{y_{i,t}^{\mathrm{BDD}}}{y_{i,0}^{\mathrm{BDD}}}\right)^{2}.$ (24)
Here $y_{i,t}^{\mathrm{NoCC}}$ is GDP per capita obtained under no climate impact by solving the following optimal growth model with a choice of $(g_{i,0},d_{i},\pi_{1,i},\pi_{2,i})$ and its associated TFP $A_{it}$:
$\max_{c_{i,t}}$ $\sum_{t=0}^{\infty}\beta^{t}u(c_{i,t})L_{i,t},$ (25)
where $y_{i,t}^{\mathrm{NoCC}}=A_{i,t}K_{i,t}^{\alpha}L_{i,t}^{-\alpha}$ and $u(c_{i,t})$ is defined as in equation (15); and $y_{i,t}$ is GDP per capita under climate impact by solving the following optimal growth model:
$\max_{c_{i,t}}$ $\sum_{t=0}^{\infty}\beta^{t}u(c_{i,t})L_{i,t},$ (26)
$\mathrm{s.t.}$ $K_{i,t+1}=(1-\delta)K_{i,t}+\left(y_{i,t}-c_{i,t}\right)L_{i,t},$